# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v4

@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [265]. SoundStream [368] is one such model, which can encode audio content at various bitrates. It is based on a residual vector quantizer (RVQ) which trades off rate, distortion and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [369] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [370], which performs hierarchical quantization at variable frame rates.
More recently, researchers have started to exploit the use of LLMs for audio compression, leveraging the audio generation/synthesis abilities of generative models. UniAudio 1.5 [371] is one of such attempts, which converts an audio into the textural space, which can be represented by a pre-trained LLM that shares a similar backbone of UniAudio [372], a universal audio foundation model. LFSC is another neural audio codec based on LLMs, which achieved fast LLM training and inference through finite scalar quantization and adversarial training.
@@ 3.7 Visual Quality Assessment
In recent years, quality assessment methods have been enhanced using deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing work in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [373, 374, 375].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference (un processed) content. These are referred to as full-reference and no-reference models [^55] Reduced-reference quality metrics do exist in the literature, but research in this field has been less active in recent years.. Prior to the AI era, conventional visual quality methods often exploited characteristics of the human vision system capturing information related to structural similarity (such as in SSIM and its variants [376, 377, 378]), distortion [379, 380, 381], and artifacts [382, 383, 384]. In many cases, the extracted features are further processed by models that simulate texture masking [385], contrast sensitivity [386], and saliency [387]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one such example [335]. When neural networks are used for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolutional neural networks were used for this, with notable examples such as DeepQA [388], LPIPS [389] and CONTRIQUE [390] for image quality assessment, and TLVQA [391], C3DVQA [392] and DeepVQA [393] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class include IQT [248], TRes [249], SaTQA [250], FastVQA [251] and RankDVQA [252]. The former has been further extended as DOVER [253] and COVER [254] when aesthetic and/or semantic aspects in the content are taken into account.
@@
More recently, inspired by the success of large language models (LLMs) [4, 394] in other machine learning tasks, these have been utilized in image and video quality assessment, demonstrating significant potential to achieve better model generalization. Q-Bench [395] is one of the first attempts that employs multimodal large language models to predict the perceptual quality of images based on prompt-driven evaluation. It queries the LLMs to provide information related to the final quality rating of the input image and the quality description. This has been further extended for video quality assessment tasks in Q-Align [396]. Other notable works include X-iqe [397] that performs the quality prompt in a multi-iteration manner focusing on both image fidelity and aesthetics. Prompt-based approaches have also been proposed for differentiating the quality difference between multiple images, such as 2AFC-LMMs [398] based on a two-alternative forced choice prompt and MAP (maximum a posteriori) estimation. Moreover, recent research works also focus on using pre-trained vision-language models, such as CLIP [263], which align better image and text modalities. Important examples in this class for image quality assessment include ZEN-IQA [399], QA-CLIP [400] and PromptIQA [401]. Similar works have also been proposed for video quality assessment, such as BVQI [402, 403] and COVER [254].
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [404], CSIQ [380], TID2013 [405], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [406], KoNViD-1K [407], YouTube UGC [408] and LIVE-VQC [409] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [410] focusing on frame rates, VSR-QAD [411] on spatial resolution (or super-resolution artifacts), BAND-2k [412] on banding artifacts and Maxwell [403]/BVI-Artifact [413] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resource requirements associated with collecting subjective data using human participants in psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.
@@
To address these issues, various proxy quality metrics have been used to label images and videos, which avoid expensive subjective tests and enable the generation of a large amounts of training material with pseudo-ground-truth quality annotations. To further improve the reliability of quality labels, instead of learning the absolute values of the quality labels, ranking-inspired training strategies have been developed, which focus on improving the monotonicity characteristics of quality. Important examples based on these weakly supervised training methodologies include RankIQA [414] and UNIQUE [415] for the image quality assessment task, and VFIPS [416] and RankDVQA [252] for video quality assessment. Moreover, different self-supervised learning approaches have also been employed, which transform quality labeling to an auxiliary task. For example, CONTRIQUE [390] learns relevant features from an unannotated image database based on the prediction of distortion types and degrees through contrastive learning. This method has been further applied to video quality assessment, resulting in a contrastive video quality estimator, CONVIQT [417]. More recently, quality-aware contrastive loss has been designed in [418, 255] to stabilize the learning process.
@@ 3.7.2 Performance and main challenges
Due to the lack of standard test conditions and limited model generalization within many existing image and video quality assessment models, deep compression methods are typically trained and benchmarked using different databases in conjunction with intra-database cross-validation. This can result in inconsistent evaluation results and conclusions. To enable a fair and meaningful comparison, various challenges and contests have been held for visual quality assessment. The Sixth Challenge on Learned Image Compression (CLIC) [336] associated with the Data Compression Conference 2024 is one of the latest examples which includes two quality assessment tracks for image and video compression. The best performer in the video quality assessment track achieves a Spearman Ranking Correlation Coefficient value of 0.825 [252], which is based on a ranking-inspired training methodology. Other notable challenges include the IEEE/CVF WACV 2023 HDR VQA Grand Challenge and the Video Super-Resolution Quality Assessment Challenge in ECCV 2024, which focus on high dynamic range and super-resolved content, respectively.
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodologies, challenges remain, including limited model generalization and high computational complexity.
@@
Another important use of quality metrics is as embedded loss functions for image and video processing optimization. This requires further capability and robustness, alongside complexity reduction, all topics to be addressed in future work.
@@ 4.1 Challenges for AI in the Creative Sector
One of the primary challenges for artists engaging with modern generative AI and LLMs is the lack of consistent, controllable outputs. These models operate via stochastic sampling from high-dimensional latent spaces, meaning that identical prompts can yield different results across runs. This unpredictability can make it challenging for artists to achieve, or iterate toward, a precise creative vision. Although prompt engineering has emerged as a technique to guide model behavior, it requires technical knowledge and iterative refinement, which may not align with the intuitive or exploratory approaches common in artistic practice.
Moreover, there can be a fundamental tension between the structured nature of current AI pipelines and current production, often improvisational, workflows used in creative disciplines. Many generative tools were originally designed for tasks like software development, content automation, or optimization [419], and are less well suited for open-ended, exploratory creation. Artists typically work in cycles of ideation, experimentation, and revision—processes that demand fluid, real-time interaction and control, which existing AI systems struggle to support. These limitations point to a gap in current AI design: a need for systems that not only generate high-quality content but also adapt to the iterative, interpretive nature of artistic production. One possible approach to address these challenges is a reinforcement of top-down creative workflows, where artists define high-level concepts, themes, or goals via text prompts before refining specific outputs. This approach helps align AI-generated results with artistic intent, offering a degree of control over inherently stochastic systems.
@@
There are also broader concerns that persist regarding the long-term impact of AI on the creative industries, particularly with the potential emergence of artificial general intelligence (AGI). Envisioned by organizations like OpenAI, DeepMind, and Anthropic, AGI could surpass human cognitive abilities, raising ethical and existential questions about the role of human agency in artistic expression.
@@ 4.2 Ethical Issues, Fakes and Bias
Diversity and representation is a key issue when using AI tools. Unified Concept Editing [54] has been proposed as a basis for image generation in digital mediums. This aims to ensure the production of safe content with diverse representation, reducing gender and racial biases. Hallucination in generative AI (the production of outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements) is a further cause of concern. These undermine trust in AI processes and can be due to limitations in the training data, biases in the model architecture or imperfections in the optimization process. Hallucinations associated with LLMs are one of the issues highlighted by the UK Government [1], alongside bias, regurgitation of private data, difficulties with multi-step tasks and challenges in interpreting black-box processes.
Governments across the world are increasingly expressing concerns about the challenges and uncertainties that generative AI technologies pose to rights holders and human creativity [2]. Generative AI presents substantial legal challenges, including the copyright status of AI-generated work and the intellectual property and copyright implications of the datasets used in training AI models. Viewpoints on this issue do however differ. For example, the track “Heart on My Sleeve,” penned by an (as yet unidentified) human author, featured AI-generated vocals that replicated the voices of Drake and The Weeknd. Released independently on April 4, 2023, it was accessible via streaming platforms including Apple Music, Spotify, and YouTube. The song quickly became viral, accumulating over 20 million views across all platforms [^58] https://www.nbcnews.com/pop-culture/viral-ai-powered-drake-weeknd-song-removed-streaming-services-rcna80098, prior to its removal by Universal Music Group, Drake’s recording label. In contrast, Canadian artist Grimes has extended an invitation to musicians to emulate her voice via AI for the creation of new musical pieces, stipulating that the lyrics should not be harmful. She has advocated for the democratization of art and the abolition of copyright [^59] https://www.bbc.com/news/entertainment-arts-65385382. Additionally, Grimes has employed AI to design visual content for her LED backdrop at Coachella in 2024.

### v5

@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [Oord:Neural:2017]. SoundStream [zeghidour2021soundstream] is one such model, which can encode audio content at various bitrates. It is based on a residual vector quantizer (RVQ) which trades off rate, distortion and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [kumar2024high] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [siuzdak2024snac], which performs hierarchical quantization at variable frame rates.
More recently, researchers have started to exploit the use of LLMs for audio compression, leveraging the audio generation/synthesis abilities of generative models. UniAudio 1.5 [yang2024uniaudio] is one of such attempts, which converts an audio into the textural space, which can be represented by a pre-trained LLM that shares a similar backbone of UniAudio [yang2023uniaudio], a universal audio foundation model. LFSC is another neural audio codec based on LLMs, which achieved fast LLM training and inference through finite scalar quantization and adversarial training.
@@ 3.7 Visual Quality Assessment
In recent years, quality assessment methods have been enhanced using deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing work in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [zhai2020perceptual, zheng2024video, zhang2024quality].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference (un processed) content. These are referred to as full-reference and no-reference models [^55] Reduced-reference quality metrics do exist in the literature, but research in this field has been less active in recent years.. Prior to the AI era, conventional visual quality methods often exploited characteristics of the human vision system capturing information related to structural similarity (such as in SSIM and its variants [Bovik_SSIM, wang2003multiscale, rehman2015display]), distortion [chandler2007vsnr, larson2010most, STMAD], and artifacts [ou2010perceptual, zhu2014no, zhang2015perception]. In many cases, the extracted features are further processed by models that simulate texture masking [helmholtz1896handbook], contrast sensitivity [kelly1977visual], and saliency [itti2001computational]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one such example [VMAFblog]. When neural networks are used for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolutional neural networks were used for this, with notable examples such as DeepQA [kim2017deep], LPIPS [zhang2018unreasonable] and CONTRIQUE [madhusudana2022image] for image quality assessment, and TLVQA [korhonen2019two], C3DVQA [xu2020c3dvqa] and DeepVQA [kim2018deep] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class include IQT [cheon2021perceptual], TRes [golestaneh2022no], SaTQA [shi2024transformer], FastVQA [wu2022fast] and RankDVQA [feng2024rankdvqa]. The former has been further extended as DOVER [wu2023exploringvideo] and COVER [he2024cover] when aesthetic and/or semantic aspects in the content are taken into account.
@@
More recently, inspired by the success of large language models (LLMs) [openai:gpt4:2023, touvron2023llama] in other machine learning tasks, these have been utilized in image and video quality assessment, demonstrating significant potential to achieve better model generalization. Q-Bench [wu2024qbench] is one of the first attempts that employs multimodal large language models to predict the perceptual quality of images based on prompt-driven evaluation. It queries the LLMs to provide information related to the final quality rating of the input image and the quality description. This has been further extended for video quality assessment tasks in Q-Align [wu2024qalign]. Other notable works include X-iqe [chen2023x] that performs the quality prompt in a multi-iteration manner focusing on both image fidelity and aesthetics. Prompt-based approaches have also been proposed for differentiating the quality difference between multiple images, such as 2AFC-LMMs [zhu20242afc] based on a two-alternative forced choice prompt and MAP (maximum a posteriori) estimation. Moreover, recent research works also focus on using pre-trained vision-language models, such as CLIP [radford2021learning], which align better image and text modalities. Important examples in this class for image quality assessment include ZEN-IQA [miyata2024zen], QA-CLIP [pan2023quality] and PromptIQA [chen2025promptiqa]. Similar works have also been proposed for video quality assessment, such as BVQI [wu2023exploring, wu2023towards] and COVER [he2024cover].
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [sheikh2006astatistical], CSIQ [larson2010most], TID2013 [ponomarenko2013color], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [seshadrinathan2010study], KoNViD-1K [hosu2017konstanz], YouTube UGC [wang2019youtube] and LIVE-VQC [sinno2018large] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [madhusudana2021subjective] focusing on frame rates, VSR-QAD [zhou2024database] on spatial resolution (or super-resolution artifacts), BAND-2k [chen2024band2k] on banding artifacts and Maxwell [wu2023towards]/BVI-Artifact [feng2024bvi] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resource requirements associated with collecting subjective data using human participants in psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.
@@
To address these issues, various proxy quality metrics have been used to label images and videos, which avoid expensive subjective tests and enable the generation of a large amounts of training material with pseudo-ground-truth quality annotations. To further improve the reliability of quality labels, instead of learning the absolute values of the quality labels, ranking-inspired training strategies have been developed, which focus on improving the monotonicity characteristics of quality. Important examples based on these weakly supervised training methodologies include RankIQA [liu2017rankiqa] and UNIQUE [zhang2021uncertainty] for the image quality assessment task, and VFIPS [hou2022perceptual] and RankDVQA [feng2024rankdvqa] for video quality assessment. Moreover, different self-supervised learning approaches have also been employed, which transform quality labeling to an auxiliary task. For example, CONTRIQUE [madhusudana2022image] learns relevant features from an unannotated image database based on the prediction of distortion types and degrees through contrastive learning. This method has been further applied to video quality assessment, resulting in a contrastive video quality estimator, CONVIQT [madhusudana2023conviqt]. More recently, quality-aware contrastive loss has been designed in [zhao2023quality, peng2024rmt] to stabilize the learning process.
@@ 3.7.2 Performance and main challenges
Due to the lack of standard test conditions and limited model generalization within many existing image and video quality assessment models, deep compression methods are typically trained and benchmarked using different databases in conjunction with intra-database cross-validation. This can result in inconsistent evaluation results and conclusions. To enable a fair and meaningful comparison, various challenges and contests have been held for visual quality assessment. The Sixth Challenge on Learned Image Compression (CLIC) [clic] associated with the Data Compression Conference 2024 is one of the latest examples which includes two quality assessment tracks for image and video compression. The best performer in the video quality assessment track achieves a Spearman Ranking Correlation Coefficient value of 0.825 [feng2024rankdvqa], which is based on a ranking-inspired training methodology. Other notable challenges include the IEEE/CVF WACV 2023 HDR VQA Grand Challenge and the Video Super-Resolution Quality Assessment Challenge in ECCV 2024, which focus on high dynamic range and super-resolved content, respectively.
@@
Although significant progress has been made in the past few years in visual quality assessment, including new models and training methodologies, challenges remain, including limited model generalization and high computational complexity. Another important use of quality metrics is as embedded loss functions for image and video processing optimization. This requires further capability and robustness, alongside complexity reduction, all topics to be addressed in future work.
3.8 Summary of AI technologies for creative industries
This section consolidates the preceding discussion by providing a comparative overview of the main classes of AI models shaping contemporary creative practice. While earlier sections examined individual technologies in detail, the following summary highlights how these models—ranging from large language models (LLMs) and diffusion models (DMs) to Neural Radiance Fields (NeRFs) and Implicit Neural Representations (INRs)—differ in application domains, key advantages, and persistent limitations. This synthesis enables readers from both creative and technical backgrounds to discern where each model type contributes most effectively, where challenges remain, and how these approaches collectively reshape workflows across the creative industries. Table 2 summarizes these relationships and serves as a conceptual reference for future research, mapping core model classes to their applications, strengths, and constraints to guide the evaluation and development of emerging AI methods in the creative industries.
Table 2: Comparative summary of key AI technologies for creative industries
Technology Core Mechanism Creative Applications Key Strengths Main Limitations Accessibility
Transformers/ Attention Self-attention mechanisms capture global dependencies Image/video restoration, super-resolution, segmentation, object detection, tracking Parallel processing, long-range context, scalable, state-of-the-art performance High memory usage, quadratic complexity, requires large training datasets High: SwinIR, SAM, Restormer (open-source)
Large Language Models (LLMs) Language understanding and generation Text generation, dialogue systems, screenwriting, storyboarding, code assistance, script analysis Highly flexible across domains; capable of reasoning and context understanding; supports creative ideation and natural language interaction Hallucinations and factual errors; limited interpretability; potential bias from training data; copyright and authorship ambiguity Moderate: GPT-4, Claude (APIs); LLaMA, Qwen (open weights)
Diffusion Models (DMs) Iterative denoising from Gaussian noise Image synthesis, Text-to-image/video generation, animation, VFX, design prototyping, restoration, style transfer, inpainting, restoration High-quality diverse outputs; stable training; handles complex distributions; photorealistic; fine-grained control through prompts or conditioning inputs Computationally expensive; slow sampling; temporal flicker in videos; difficulty ensuring semantic or style consistency High: Stable Diffusion (open); DALL·E 3, Sora (APIs)
Neural Radiance Fields (NeRFs) Volumetric scene representation via MLPs 3D scene reconstruction, virtual production, spatial storytelling, AR/VR content creation Photorealistic rendering from sparse views; compact scene representation; supports dynamic viewpoint changes Limited to static or semi-static scenes; slow training; poor performance in textureless regions; large memory footprint Moderate: Instant-NGP, Mip-NeRF (open); requires CV expertise
Implicit Neural Representations (INRs) Coordinate-based continuous functions Video/image compression; super-resolution; 3D reconstruction; dynamic scene modeling Continuous scene representation; parameter-efficient encoding; smooth interpolation and compact storage Challenging to generalize across scenes; training instability; limited semantic control Moderate: NeRV, COIN (open); technical expertise required
Gaussian Splatting Explicit 3D Gaussians for scene representation Real-time novel view synthesis; 3D scene reconstruction; VR/AR;interactive rendering Fast training (minutes) and rendering (real-time); explicit representation; easy manipulation Lower quality than NeRF for complex scenes, requires SfM initialization, memory intensive High: 3D-GS, 4D-GS (open); LUMA AI, Polycam (free tools)
* Product and platform data (e.g., Sora, Gemini, Stable Diffusion 3) are accurate as of mid-2025 and may evolve rapidly.
@@ 4.1 Challenges for AI in the Creative Sector
Artists are already exploring how to bridge structured nature of current creative AI and more traditional (analog) workflows. This includes using multiple methods, models, or tools to create new works. For example, artists can use tools to iterate on existing or past works; upload and fuse analog works or output from different Generative AI tools to further intervene in the AI generation process; and composite outputs to reconstruct or fill in missing or damaged parts of a work. However, one of the primary challenges for artists engaging with modern generative AI and LLMs is the lack of consistent, controllable outputs. These models operate via stochastic sampling from high-dimensional latent spaces, meaning that identical prompts can yield different results across runs. This unpredictability can make it challenging for artists to achieve, or iterate toward, a precise creative vision. Although prompt engineering has emerged as a technique to guide model behavior, it requires technical knowledge and iterative refinement, which may not align with the intuitive or exploratory approaches common in artistic practice.
Moreover, there can be a fundamental tension between the structured nature of current AI pipelines and current production, often improvisational, workflows used in creative disciplines. Many generative tools were originally designed for tasks like software development, content automation, or optimization [zhong:LDB:2024], and are less well suited for open-ended, exploratory creation. Artists typically work in cycles of ideation, experimentation, and revision—processes that demand fluid, real-time interaction and control, which existing AI systems struggle to support. These limitations point to a gap in current AI design: a need for systems that not only generate high-quality content but also adapt to the iterative, interpretive nature of artistic production. One possible approach to address these challenges is a reinforcement of top-down creative workflows, where artists define high-level concepts, themes, or goals via text prompts before refining specific outputs. This approach helps align AI-generated results with artistic intent, offering a degree of control over inherently stochastic systems.
A further issue concerns creative authorship and ownership. As AI systems increasingly contribute to the ideation and execution of creative work, the line between human and machine authorship becomes blurred. Determining where creative credit lies—whether with the prompt designer, the model developer, or the AI system itself—poses significant legal and ethical challenges. Moreover, generative systems often reproduce stylistic elements from training data, raising questions about originality and cultural appropriation in AI-assisted creation.
@@
There are also broader concerns that persist regarding the long-term impact of AI on the creative industries, economies, and labor markets. As automation accelerates, socioeconomic disparities may widen between those who can afford access to powerful generative systems and those who cannot. Freelancers and smaller studios risk being marginalized by large organizations with access to proprietary datasets and substantial compute resources. The potential emergence of artificial general intelligence (AGI). Envisioned by organizations like OpenAI, DeepMind, and Anthropic, AGI could surpass human cognitive abilities, raising ethical and existential questions about the role of human agency in artistic expression. Ensuring equitable access to AI tools and fair distribution of creative value will therefore be crucial to sustaining diversity, inclusion, and innovation within the sector.
@@ 4.2 Ethical Issues, Fakes and Bias
Diversity and representation is a key issue when using AI tools. Unified Concept Editing [Gandikota:Unified:2024] has been proposed as a basis for image generation in digital mediums. This aims to ensure the production of safe content with diverse representation, reducing gender and racial biases. Hallucination in generative AI (the production of outputs that are not faithful representations of reality but instead contain imagined or unrealistic elements) is a further cause of concern. These undermine trust in AI processes and can be due to limitations in the training data, biases in the model architecture or imperfections in the optimization process. Hallucinations associated with LLMs are one of the issues highlighted by the UK Government [UK:Large:2024], alongside bias, regurgitation of private data, difficulties with multi-step tasks and challenges in interpreting black-box processes.
Governments across the world are increasingly expressing concerns about the challenges and uncertainties that generative AI technologies pose to rights holders and human creativity [Jeary2024]. Generative AI presents substantial legal challenges, including the copyright status of AI-generated work and the intellectual property and copyright implications of the datasets used in training AI models. Viewpoints on this issue do however differ. For example, the track “Heart on My Sleeve,” penned by an (as yet unidentified) human author, featured AI-generated vocals that replicated the voices of Drake and The Weeknd. Released independently on April 4, 2023, it was accessible via streaming platforms including Apple Music, Spotify, and YouTube. The song quickly became viral, accumulating over 20 million views across all platforms [^58] https://www.nbcnews.com/pop-culture/viral-ai-powered-drake-weeknd-song-removed-streaming-services-rcna80098, prior to its removal by Universal Music Group, Drake’s recording label. In contrast, Canadian artist Grimes has extended an invitation to musicians to emulate her voice via AI for the creation of new musical pieces, stipulating that the lyrics should not be harmful. She has advocated for the democratization of art and the abolition of copyright [^59] https://www.bbc.com/news/entertainment-arts-65385382. Additionally, Grimes has employed AI to design visual content for her LED backdrop at Coachella in 2024.